OpenStudy (anonymous):

Chocolates costing $8 per pound are to be mixed with chocolates costing $3 per pound to make a 20 pound mixture. If the mixture is to sell for $5 per pound, how many pounds of each chocolate should be used? Which of the following equations could be used to solve the problem? 8x + 3x = 5(20) 8x + 3(20) = 5(x + 20) 8x + 3(20 - x) = 5(20)

OpenStudy (anonymous):

I think B? @parthkohli

OpenStudy (anonymous):

haha thanks.

OpenStudy (anonymous):

Or! maybe A because I don't think you'd multiply the 3 by 20 .-. lol Parth will help us

OpenStudy (anonymous):

okay, im so confused by this stuff lol

OpenStudy (anonymous):

\[8x+3(20-x)=5\times 20\] is your equation

OpenStudy (anonymous):

Satellite is here...I will shut up now.

OpenStudy (anonymous):

thanks ya'll

OpenStudy (anonymous):

we reason as follows: if you put the amount of $8 a pound chocolate as say \(x\) then since the total number of pounds is 20, the amount of $3 a pound chocolate must be \(20-x\)

OpenStudy (anonymous):

\(x\) pounds at $8 per pound costs \(8x\) and \(20-x\) pounds at $3 costs \(3\times (20-x)\) for a total cost of \[8x+3(20-x)\]

OpenStudy (anonymous):

wow your really smart

OpenStudy (anonymous):

since you know the total cost is $5 times 20 = 100 you can set \[8x+3(20-x)=100\] and solve for \(x\)